OpenStudy (anonymous):

fration divison 1/(2/3)=3/2?

hero (hero):

\[1 \div \frac{2}{3} = 1 \times \frac{3}{2} = \frac{3}{2}\] Correct

OpenStudy (anonymous):

so its bigger than one?

hero (hero):

Yes \[\frac{3}{2} = \frac{2 + 1}{2} = \frac{2}{2} + \frac{1}{2} = 1 \frac{1}{2}\]

OpenStudy (anonymous):

thank you!